OpenStudy (anonymous):

Help Please The sum of the angle measure of a polygon with s sides is 2,520. Find s a.)14 b.)16 c.)18 d.)15

OpenStudy (solomonzelman):

\(\large\color{black}{ (s-2)(180)=2,520 }\)

OpenStudy (solomonzelman):

solve for s

OpenStudy (anonymous):

-360=2,520 i think i did that totally wrong

OpenStudy (solomonzelman):

do you know why I wrote \(\large\color{black}{ (s-2)(180)=2,520 }\) ?

OpenStudy (anonymous):

no why??

OpenStudy (solomonzelman):

The sum of angle measures in a s-gon, with number sides s, is given by \(\large\color{black}{ (s-2)(180)}\) We know that in s-gon the sum of measures is equal to \(\large\color{black}{ 2,520 }\) So, we can write that the sum of the angles in an s-gon is equivalent to 2520. equation for this is, \(\large\color{black}{ (s-2)(180)=2,520 }\)

OpenStudy (solomonzelman):

Makes sense?

OpenStudy (anonymous):

okay but im not understanding how i solve that.. :(

OpenStudy (solomonzelman):

okay, you have a product of (s-2) and 180 is equal to 2520. First, isolate the (s-2) by dividing both sides by 180. 2520 divided by 180 is 14 and "(s-2)180" divided by 180 is just (s-2). So you get \(\large\color{black}{ s-2=14 }\) And the last step, you can do on your own:)

OpenStudy (anonymous):

so its 16 :) Thank you so much :)

OpenStudy (solomonzelman):

No problem.